OpenStudy (anonymous):

So how do I solve absolute value problems? EXAMPLES: 6 + [4], 8-[3+5], [9-7+3], [-4] + [-5] I NEED TO KNOW HOW TO DO ALL DIFFERENT TYPES

OpenStudy (anonymous):

absolute value takes any number and makes it positive, so [x]=x and [-x]=x

OpenStudy (anonymous):

http://www.purplemath.com/modules/absolute.htm

OpenStudy (anonymous):

Wait, those brackets mean absolute value? I thought they were just brackets.

OpenStudy (amistre64):

do the equations inside the |...| and whatever number you get; drop the sign

OpenStudy (amistre64):

some people cant find the "|" key :) on my keyboard its right above the "enter"

OpenStudy (anonymous):

On mine too. I just don't understand how you assume the [] is an alternative for absolute value symbols ||

OpenStudy (anonymous):

because he said they were absolute value problems

OpenStudy (amistre64):

maybe its a british thing :)

OpenStudy (anonymous):

Okay, I didn't read the first line. I'm not british.

OpenStudy (anonymous):

So @amistre64, your saying i'd start with 8-[3+5] and then.... 8-[8]????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

[-4] + [-5] = ....isn't it the same?

OpenStudy (anonymous):

yes, so what do you get, I just want to make sure you understand how to solve them that is all

OpenStudy (amistre64):

8-|8| = 8-8 = 0 yes |-4| + |-5| = 4+5 = 9

OpenStudy (anonymous):

ooooooooh! =) Thanks amistre64!!!!

OpenStudy (amistre64):

all absolute value are is a distance, a length, an absolute value. it doesnt care what teh sign directions are :)

OpenStudy (anonymous):

solve everything in the absolute value functions first, then make whatever that is positive (as amistre says), then solve the rest of the equation... similar to parenthesis, do em first, then the rest...

OpenStudy (amistre64):

a length of 4 add to a length of 5 = a length of 9

OpenStudy (anonymous):

Just treat absolute values like other brackets or parens, but before removing them you have to make whatever is inside them positive.

OpenStudy (anonymous):

The only time they get 'tricky' is when solving for a variable inside the absolute values.

OpenStudy (anonymous):

^whether that length of four goes left or right from 0 on the number line, its still a length of 4, thats what absolute values are

OpenStudy (anonymous):

can someone please tell me why my open study isnt working right when i log in the column on the lefy is alwayss loading slow and i can not get into the chat